# Compliance Brain Assistant: Conversational Agentic AI for Assisting Compliance Tasks in Enterprise Environments

### v2

@@ 2. Architecture
Router. At the core of the new architecture is a routing-based framework that intelligently directs user queries to the appropriate processing track. The Router leverages a potent LLM model, GPT-4o (Hurst et al., 2024), to determine whether a query requires the full agentic workflow or can be more efficiently handled by the FastTrack flow. This decision-making process is important for reducing latency by avoiding unnecessary steps for queries that do not require complex reasoning or multi-step tool use.
@@ 3.1.2. Test Methodology
The router was presented with the 15 queries in isolation (without prior conversation context) and forced to choose between FastTrack or FullAgentic. We compared its decisions against the expert labels using standard classification metrics.
@@ 3.1.3. Results
Table 1. Router accuracy results: (a) confusion matrix and (b) classification metrics.
[…]
FastTrack 7 1
FullAgentic 1 6
@@
Class Recall Precision
FastTrack 87.5% 87.5%
FullAgentic 85.7% 85.7%
@@
The router demonstrates strong performance (86.7% overall accuracy) in distinguishing between generic and context-rich queries, validating its role as the decision-making backbone of CBA. The errors mostly arise from borderline cases (e.g., queries that can be considered to only marginally benefit from having artifact support), and reveal clear improvement opportunities - we have added these misclassified examples to our in-context dataset and will retrain the router periodically using this growing collection of edge cases. This iterative refinement process will enable continuous improvement in routing accuracy while maintaining low latency through the lightweight classifier architecture. Future work will explore incorporating real-time feedback from privacy specialists to further enhance decision quality.
@@ 3.2. End-to-end Performance
    Routing-based: This condition used CBA with the Router module as previously described, which dynamically selects the most suitable workflow (FastTrack or FullAgentic) based on the input prompt.
@@ 3.2.3. Results
The experimental results are presented in Tables 2, 3 and 3 across different benchmarks. First, CBA substantially outperforms the vanilla LLM across all quality metrics on all the benchmarks, showing that our compliance-oriented enhancements are effective (e.g., on Compliance Knowledge Benchmark, the best PBA configuration more than doubles the average match rate (83.7% vs. 41.7%) and the LLM-judge’s pass rate (82.0% vs. 20.0%) compared to the vanilla LLM). In fact, the Router mode leads in terms of quality metrics, proving to be the most performant mode on all benchmarks except the the artifact understanding one (which we find reasonable as this benchmark requires agentic actions for all samples). From the latency perspective, we notice that because the router needs one additional LLM call to determine the appropriate mode to engage for the given user query, it tends to incur longer time compared to directly entering specific modes. In addition, it is largely expected that FullAgentic on average incurs higher overhead because agent performs multi-step actions which is intrinsically time-consuming; we do note that FullAgentic somehow runs faster than other setups on Compliance Knowledge Benchmark, which after inspecting logs we find is due to that FullAgentic in cases chooses incorrect but fast tools (e.g., keyword-based knowledge retrieval), resulting in low latency but also low quality as shown in Tables 2. Overall, the experimental results suggest that using the router is generally a robust and reliable choice for achieving both high response quality, and acceptable response latency at the same time, across a range of tasks.

### v3

@@ 2. Architecture
Router. At the core of the new architecture is a routing-based framework that intelligently directs user queries to the appropriate processing track. The Router leverages an efficient while effective LLM model, Llama3.1-8B, to determine whether a query requires the full agentic workflow or can be more efficiently handled by the FastTrack flow. This decision-making process is important for reducing latency by avoiding unnecessary steps for queries that do not require complex reasoning or multi-step tool use.
@@ 3.1.2. Test Methodology
The router was presented with the 15 queries in isolation (without prior conversation context) and forced to choose between FastTrack or FullAgentic. We compared its decisions against the expert labels using standard classification metrics. We tested different sizes of Llama 3 series models (3.1-8B, 3.2-1B, 3.2-3B, 3.3-70B) to compare their efficiency and effectiveness.
@@ 3.1.3. Results
Table 1. Router accuracy and efficiency results: (a) confusion matrix, (b) classification metrics and (c) latency
Actual/Predicted 1B 3B 8B 70B
FastTrack/FastTrack 4 7 7 7
FastTrack/FullAgentic 4 1 1 1
FullAgentic/FastTrack 2 1 1 1
FullAgentic/FullAgentic 5 6 6 6
@@
Class 1B 3B 8B 70B
FastTrack Precision 66.7% 87.5% 87.5% 87.5%
FastTrack Recall 50.0% 87.5% 87.5% 87.5%
FullAgentic Precision 55.6% 85.7% 85.7% 85.7%
FullAgentic Recall 71.4% 85.7% 85.7% 85.7%
@@
Model 1B 3B 8B 70B
Latency (s) 0.77 0.88 1.61 3.15
(c) Average Latency per Request
The results are presented in Table 1. Given the relatively straightforward nature of this classification task for LLMs, any Llama model above 3B parameters performs effectively. Among these, the router demonstrates strong performance (86.7% overall accuracy) in distinguishing between generic and context-rich queries, validating its role as the decision-making backbone of CBA. The errors mostly arise from borderline cases (e.g., queries that can be considered to only marginally benefit from having artifact support), and reveal clear improvement opportunities - we have added these misclassified examples to our in-context dataset and will retrain the router periodically using this growing collection of edge cases. This iterative refinement process will enable continuous improvement in routing accuracy while maintaining low latency through the lightweight classifier architecture. We have selected the Llama3.1-8B model as the base for the router due to its strong performance, sufficient speed for our use case, and greater potential for future enhancements compared to the smaller 3B model. Future work will explore incorporating real-time feedback from privacy specialists to further enhance decision quality.
@@ 3.2. End-to-end Performance
    Routing-based: This condition used CBA with the Router module based on Llama3.1-8B as previously described, which dynamically selects the most suitable workflow (FastTrack or FullAgentic) based on the input prompt.
@@ 3.2.3. Results
The experimental results are presented in Tables 2, 3 and 4 across different benchmarks. First, CBA substantially outperforms the vanilla LLM across all quality metrics on all the benchmarks, showing that our compliance-oriented enhancements are effective (e.g., on Compliance Knowledge Benchmark, the best CBA configuration more than doubles the average match rate (83.7% vs. 41.7%) and the LLM-judge’s pass rate (82.0% vs. 20.0%) compared to the vanilla LLM). In fact, the Router mode leads in terms of quality metrics, proving to be the most performant mode on all benchmarks except the the artifact understanding one (which we find reasonable as this benchmark requires agentic actions for all samples). From the latency perspective, we notice that because the router needs one additional LLM call to determine the appropriate mode to engage for the given user query, it tends to incur longer time compared to directly entering specific modes. In addition, it is largely expected that FullAgentic on average incurs higher overhead because agent performs multi-step actions which is intrinsically time-consuming, especially for datasets with difficult questions like the Regulation Knowledge Benchmark (Table 3). For Compliance Knowledge Benchmark (Table 2), FullAgentic runs surprisingly faster than other setups. After inspecting logs, we found that FullAgentic chose incorrect but fast tools (e.g., keyword-based knowledge retrieval) and generated the answer with only one step, resulting in low latency but also low quality. On the other hand, the questions from this particular dataset (understanding compliance concepts) can be better answered with RAG with a large knowledge base. That’s why FastTrack and router are slightly slower but perform better on this dataset. Overall, the experimental results suggest that using the router is generally a robust and reliable choice for achieving both high response quality, and acceptable response latency at the same time, across a range of tasks.